OpenStudy (adamaero):

Is the integral of ln(0) an improper integral?

OpenStudy (adamaero):

Since taking the limit of it, you would be approaching \[-\infty\]

OpenStudy (adamaero):

@jdoe0001 @e.mccormick @Compassionate @kropot72 @uri @timo86m @Ashleyisakitty

OpenStudy (adamaero):

@Preetha @bibby @karatechopper @Isaiah.Feynman @poopsiedoodle @camerondoherty @superhelp101

OpenStudy (poopsiedoodle):

If it helps, I just put ln(0) in my calculator and got a domain error. I don't know anything about integrals though :/

OpenStudy (adamaero):

ya, that didn't help. thanks though

OpenStudy (zarkon):

your question, as stated, makes little sense. Try giving us the original question.

OpenStudy (adamaero):

\[\int\limits_{1}^{2} \ln(x-1) dx\]

OpenStudy (adamaero):

I meant evaluated at ln(1-1)

OpenStudy (zarkon):

the integral you typed is an improper integral